OpenStudy (anonymous):

show that if f is odd and find the inverse of f f(x)= ln(x+sqrt(x^2+1))

OpenStudy (anonymous):

f is odd if f(-x)=-f(x)

OpenStudy (anonymous):

@ mukushl how?????

OpenStudy (anonymous):

\[f(-x)=\ln (-x+\sqrt{1+x^2)}\] and \[-f(x)=-\ln (x+\sqrt{1+x^2)}\]

OpenStudy (anonymous):

\[ f(x)= \sinh^{-1}(x) \] and sinh is odd. The inverse of f is sinh(x)

OpenStudy (anonymous):

See http://en.wikipedia.org/wiki/Hyperbolic_function

OpenStudy (anonymous):

eliassaab's way is much better

OpenStudy (anonymous):

@mukushla f(-x) not equal -f(x)

OpenStudy (anonymous):

so f(x) is not odd

OpenStudy (anonymous):

But in question it is written SHOW THAT so f(x) is an odd function..

OpenStudy (anonymous):

This is what I want

OpenStudy (anonymous):

@waterineyes is f an odd function?

OpenStudy (anonymous):

im wondering because it seems f(-x) not equal -f(x)

OpenStudy (anonymous):

Question says f is an odd function we have to just show it and I am unable in doing that.. Do you think that ln(......) is any formula or not??

OpenStudy (anonymous):

im sorry \[\frac{1}{x+\sqrt{1+x^2}}=-x+\sqrt{1+x^2}\] then f is odd

OpenStudy (anonymous):

Guys, f(x) is odd. See my post above and copy and paste in your browser to see that. http://www.wolframalpha.com/input/?i=Plot [Log[x+%2B+Sqrt[1+%2B+x^2]]%2C+{x%2C+-100%2C+100}]

OpenStudy (anonymous):

That is what I was thinking...

OpenStudy (anonymous):

I think openstudy should change the way medals are awarded. The only one who got a medal here is one who did not contribute to the solution.

OpenStudy (anonymous):

for the inverse: \[y=\ln(x+\sqrt{x^2+1})\] inverse is given by: \[x=\ln(y+\sqrt{y^2+1})\] \[e^x=e^{\ln(y+\sqrt{y^2+1})}\] \[e^x=y+\sqrt{y^2+1}\] \[e^x/y=1+\sqrt{(y^2+1)/y^2}\] \[e^x/y=1+\sqrt{1+1/y^2}\] \[e^x/y-1=\sqrt{1+1/y^2}\] \[e^{2x}/y^2-2e^x/y+1=1/y^2+1\] \[e^{2x}-2e^xy+y^2=1+y^2\] \[y=(e^{2x}-1)/2e^x\] <-answer

OpenStudy (anonymous):

but icant find perfect answer for f is odd

OpenStudy (anonymous):

You want to prove f is odd??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you know what is an odd function??

OpenStudy (anonymous):

f(-x) = -f(x)

OpenStudy (anonymous):

Yes you got it very right.. How we find first f(-x) by replacing x by (-x) Okay??

OpenStudy (anonymous):

\[f(−x)= \ln(-x + \sqrt{1 + x^2})\] Got or not??

OpenStudy (anonymous):

And now we find -f(x): \[-f(x) = -\ln(x + \sqrt{1 + x^2})\] Got till here?

OpenStudy (anonymous):

What you think @obada these two expressions are equal or not??

OpenStudy (anonymous):

not equal

OpenStudy (anonymous):

I will prove they are equal just be cool... See, here we need to solve more so that we can prove they are equal.. You know one of the properties of Log is: \[a.\ln(x ) = \ln(x)^a\] This is called Power Rule...

OpenStudy (anonymous):

im waiting you waterineyes

OpenStudy (anonymous):

\[(-1)\ln(x + \sqrt{1 + x^2}) = \ln(x + \sqrt{1 + x^2})^{-1}\] \[\ln(\frac{1}{x + \sqrt{1 + x^2}})\] This is because: \[a^{-1} = \frac{1}{a}\] So, Rationalizing the denominator we get: \[= \ln(\frac{1}{x + \sqrt{1 + x^2}} \times \frac{x - \sqrt{1 + x^2}}{x -\sqrt{1 + x^2}})\] \[= \ln(\frac{x - \sqrt{1 + x^2}}{x^2 - 1 - x^2}) = \ln(\frac{x - \sqrt{1 + x^2}}{-1})\] \[ = \ln(-x + \sqrt{1 + x^2)}\] (Just taken the -common from Numerator and this - cancels with denominator..)

OpenStudy (anonymous):

Now what you think the result matches with f(-x) or not???

OpenStudy (anonymous):

thank you very much @waterineyes just you answer me yes now iknow how is odd ? thx again

OpenStudy (anonymous):

No need to say thanks.. Welcome dear...

OpenStudy (anonymous):

@eliassaab, do you want to say anything else??

OpenStudy (anonymous):

I do not know why are we still discussing it. I gave a one line proof that f is odd at the very beginning and what is the inverse of it.

OpenStudy (anonymous):

Well, math can be solved in more than one way.. You gave your one.. so,, we gave our one...